Call Option You have taken a long position in a call option on UBR common stock. The option has an exercise price of $142 and IBM’s stock currently trades at $145. The option premium is $6 per contract.

a. What is your net profit on the option if UBR’s stock price increases to $150 at expiration of the option and you exercise the option?

b. How much of the option premium you paid is due to intrinsic value and how much due to time value? c. What is your net profit on the option if UBR’s stock price increases to $148?
Answer #1

a

Profit = exercise price-strike price-premium = 150-142-6=2

b

Intrinsic value = current price-strike price = 145-142=3

Time value = premium-Intrinsic value = 6-3=3

c

Profit = exercise price-strike price-premium = 148-142-6=0